OpenStudy (anonymous):

Find an equation in standard form for the hyperbola with vertices at (0, ±2) and foci at (0, ±11)

OpenStudy (tylerd):

alright then first thing to take note of is the vertices are at (0, 2) and 0, -2)

OpenStudy (anonymous):

Okay, got that. totally lost after that

OpenStudy (tylerd):

so \[\frac{ y^2 }{ a^2 } - \frac{ x^2 }{ b^2 } = 1\]

OpenStudy (anonymous):

okay whats the next step?? sorry im so confused with this

OpenStudy (tylerd):

our "a" = 2 because the major axis is the one with the vertices.

OpenStudy (anonymous):

so the b will be -2

OpenStudy (tylerd):

dunno that yet

OpenStudy (tylerd):

from point (0,2) to point (0,-2) = 2a

OpenStudy (tylerd):

the total length is 4, so a =2

OpenStudy (anonymous):

okay got it

OpenStudy (tylerd):

the foci are at (0,c) and (0,-c)

OpenStudy (tylerd):

the foci and the vertices share the same x value in the (x,y) because the hyperbola is vertical

OpenStudy (tylerd):

so c = 11 or (0,c)

OpenStudy (tylerd):

\[c^2=a^2+b^2\]

OpenStudy (tylerd):

\[11^2=2^2+b^2\]

OpenStudy (anonymous):

ohhhh

OpenStudy (anonymous):

i think im getting this now

OpenStudy (tylerd):

hold on though, gonna check this make sure im right first

OpenStudy (anonymous):

okay

OpenStudy (tylerd):

ya

OpenStudy (anonymous):

ok i think im getting the hang of it, im gonna show you the answer choices i have

OpenStudy (anonymous):

y squared over 4 minus x squared over 117 equals 1 y squared over 4 minus x squared over 121 equals 1 y squared over 121 minus x squared over 4 equals 1 y squared over 117 minus x squared over 4 equals 1

OpenStudy (tylerd):

\[\frac{ y^2 }{ a^2 } - \frac{ x^2 }{ b^2 } = 1\] plug the values in for a and b

OpenStudy (tylerd):

\[\frac{ y^2 }{ 4 } - \frac{ x^2 }{ ? } = 1\]

OpenStudy (tylerd):

\[11^2=2^2+b^2\]

OpenStudy (anonymous):

121?

OpenStudy (tylerd):

\[121 = 4 + b^2\]

OpenStudy (tylerd):

subtract the 4 first to isolate b^2

OpenStudy (anonymous):

117

OpenStudy (tylerd):

yep

OpenStudy (anonymous):

so y squared over 4 minus x squared over 117 equals 1

OpenStudy (tylerd):

but remember in this case, the major axis was the vertical, if its the horizontal its done differently.

OpenStudy (tylerd):

but ya thats right

OpenStudy (anonymous):

thanks man! I think i gave you a medal, im a newbie to this site

OpenStudy (tylerd):

yep, good luck